# Decoding Open-Ended Information Seeking Goals from Eye Movements in Reading

### v2

@@ Abstract
When reading, we often have specific information that interests us in a text. For example, you might be reading this paper because you are curious about LLMs for eye movements in reading, the experimental design, or perhaps you wonder “This sounds like science fiction. Does it actually work?”. More broadly, in daily life, people approach texts with any number of text-specific goals that guide their reading behavior. In this work, we ask, for the first time, whether open-ended reading goals can be automatically decoded solely from eye movements in reading. To address this question, we introduce goal decoding tasks and evaluation frameworks using large-scale eye tracking for reading data in English with hundreds of text-specific information seeking tasks. We develop and compare several discriminative and generative multimodal text and eye movements LLMs for these tasks. Our experiments show considerable success on the task of selecting the correct goal among several options, and even progress towards free-form textual reconstruction of the precise goal formulation. These results open the door for further scientific investigation of goal driven reading, as well as the development of educational and assistive technologies that will rely on real-time decoding of reader goals from their eye movements. [^1] Code is available anonymously here.
@@ 1 Introduction
Eye movements in reading are a key methodology for studying the cognitive basis of reading and language processing. When we read, our eyes move over the text in a saccadic manner, with prolonged periods of time in which the gaze is stable, called fixations, and fast transitions between fixations called saccades (Schotter & Dillon, 2025). This trajectory contains rich behavioral traces of the ways in which readers interact with texts and process language in real time (Rayner, 1998; Just & Carpenter, 1980). Understanding the nature and strength of the relation between eye movements, the text, and online language processing has been a major research avenue in the psychology of reading in the past few decades, and in recent years has also been gaining increasing interest in NLP and ML (Barrett & Hollenstein, 2020; Reich et al., 2025).
In this work, we investigate a reading scenario that is prevalent in daily life but remains understudied – seeking specific information in a text. This scenario deviates from a common implicit assumption in psycholinguistics, according to which the comprehender’s goal is constant across communicative contexts: a general understanding of the linguistic input. In practice, readers approach texts with a variety of goals and information seeking needs. Each such goal can have a profound impact on the cognitive processes of language comprehension and the corresponding behavioral traces of eye movements over the text (Radach & Kennedy, 2004; Kaakinen & Hyönä, 2010; Hahn & Keller, 2023; Shubi & Berzak, 2023).
@@
Tasks: We present a new cognitive state decoding challenge from eye movements in reading: decoding arbitrary information seeking goals over the content of specific texts. We instantiate this challenge as a goal selection task given several possible goals, and as an even more challenging open-ended goal reconstruction task where the information-seeking goal has to be generated.
Models: We introduce two strong baseline models and adjust two state-of-the-art discriminative models that combine text with eye movements for the goal selection task. We further develop two new generative text and eye movements multimodal LLMs that can be used for both goal selection and goal reconstruction.
Evaluations: We perform systematic evaluations of selection accuracy against informative baseline models at two levels of task difficulty and utilize several generation quality measures for the reconstruction task. For both tasks, we examine model generalization across new texts, new readers, and the combination of both.
@@ 2 Eye Tracking Data
We use OneStop (Berzak et al., 2025), a public eye tracking dataset collected with a high-end EyeLink 1000 Plus eye tracker. It comprises 360 adult native English speakers reading Guardian articles from the OneStopQA multiple choice reading comprehension dataset (Berzak et al., 2020; Vajjala & Lučić, 2018). In each experimental trial, a participant reads a paragraph on a screen and, on the following screen, answers a multiple-choice reading comprehension question, without the ability to return to the paragraph. 180 participants read in an information seeking condition, where they receive the question (but not the answers) before reading the paragraph, and therefore have a specific reading goal for the upcoming paragraph. The remaining 180 participants do not receive the question ahead of time, and thus have to be ready to answer any subsequent question. Here, we use the information seeking part of OneStop, which is uniquely suited for our study.
Figure 1: Example of a text from OneStop, consisting of a paragraph and three questions. Two of the questions have the same critical span: the paragraph segment essential for answering the question.
For each paragraph, a participant receives one of three possible questions over the paragraph. Each such question has a manually annotated critical span in the paragraph, which contains the information that is essential for answering the question correctly. Two of the questions share the same critical span, while the third question has a distinct critical span. This design supports a question selection task with two tiers of difficulty, distinguishing between questions with different critical spans and the more challenging variant of distinguishing between questions that share the same critical span. Figure˜1 presents an example of a paragraph, its three questions, and their corresponding critical span annotations.
The data has 486 questions over 162 paragraphs, with 20 participants answering each question. The mean question length is 10 words. The mean paragraph length is 109 words (5.8 sentences), of which the critical span is 32 words. Overall, there are 1,055,429 paragraph word tokens for which eye tracking data was collected in the information seeking regime.
@@ 3 Problem Formulation
The general problem we address is using eye movements over a text (in our setting, a text is always a paragraph) to predict which question a participant received prior to reading that text. Each text $T$ has three text specific questions, $Q_{T}=\left\{q_{1,c_{1}},q_{2,c_{2}},q_{3,c_{2}}\right\}$. The critical information for answering $q_{1,c_{1}}$ is located in the text in span $c_{1}$, while for both $q_{2,c_{2}}$ and $q_{3,c_{2}}$ it is in span $c_{2}$. In each experiment trial, a participant $P$ is presented with one question, $q_{P}\in Q_{T}$, before reading the text. The recording of the participant’s eye movements while reading the text is denoted by $E_{P,T}$.
We propose two task variants as follows.
Question Selection In the selection task, a classifier $h$ is trained to select which question from $Q_{T}$ the participant received, from their eye movement recording over the text:
$h(T,Q_{T},E_{P,T})\longrightarrow\hat{q}_{P}\in Q_{T}$
Question Reconstruction In the free-form reconstruction task, a generative model $g$ is trained to generate the question that the participant received from their eye movement recording over the text:
@@ 4 Models
Figure 2: Two model types for decoding the question presented to a participant before reading the text, from their eye movements while reading the text. (a) Discriminative models that score candidate questions and are evaluated on question selection accuracy. (b) Generative models trained to reconstruct the question presented to the reader, and evaluated on question selection accuracy (via log-likelihood assignment to each candidate question) and free-form question reconstruction quality.
@@ 4.1 Discriminative Models
We develop two simple baseline models which build on the following observations from prior literature: (i) readers tend to spend more time on question-relevant than on question-irrelevant portions of the text (Hahn & Keller, 2023; Malmaud et al., 2020; Shubi & Berzak, 2023), and (ii) question-relevant text segments tend to have higher semantic similarity to the question compared to question-irrelevant segments (Mitra et al., 2016).
@@
Question Similarity to RT-Weighted Passage (Figure˜3(a)) For a participant reading a text of $N$ words, we compute a text embedding, $\text{emb}(T)$, as a weighted average of its contextualized word embeddings, where the weights are speed-normalized word reading times, $RT(w_{i})$. Formally: $\text{emb}(T)=\sum_{i=1}^{N}\text{RT}(w_{i})\,\text{emb}(w_{i})$. The embeddings for word $w_{i}$, $\text{emb}(w_{i})$, are extracted from RoBERTa (Liu et al., 2019), and $\text{RT}(w_{i})$ is the word’s Total Fixation Duration (i.e., sum of all the fixation durations) normalized by the text’s total reading time. For each candidate question $q_{j}$, we use its [CLS] token embedding as $\text{emb}(q_{j})$, and compute its cosine similarity with $\text{emb}(T)$. We then predict the question with the highest similarity:
${\arg\max}_{j\in\{1,_{\cdots},|Q_{T}|\}}\cos(\text{emb}(T),\text{emb}(q_{j}))$
RT Similarity to Question-Word Similarities (Figure˜3(b)) An alternative to the model above. Compares the vector of word reading times for the text to the vector of word–question similarities. Specifically, for each candidate question $q_{j}$ and text word $w_{i}$, we first compute $\text{sim}(q_{j},w_{i})=\cos\bigl(\text{emb}(q_{j}),\text{emb}(w_{i})\bigr)$, where as before, $\text{emb}(w_{i})$ is a contextual word embedding from RoBERTa, and $\text{emb}(q_{j})$ is the question [CLS] embedding, yielding a length-$N$ vector of similarities $\text{sim}(q_{j},T)=[\text{sim}(q_{j},w_{1}),\text{sim}(q_{j},w_{2}),...,\text{sim}(q_{j},w_{n})]$. We select the question maximizing the cosine similarity between this vector and $\text{RT}(T)$, the length-$N$ vector of speed-normalized Total Fixation Durations for the text words:
${\arg\max}_{j\in\{1,_{\cdots},|Q_{T}|\}}\cos\bigl(\text{sim}(q_{j},T),\,\text{RT}(T)\bigr)$
@@
We further adapt two state-of-the-art encoder models for eye movements in reading to our task. Building on the approach used in Radford et al. (2019) for encoding multiple-choice reading comprehension items, we use several copies of each text, where instead of combining each copy with a candidate answer, we combine it with a candidate question, and further provide the eye movements over the text $\langle T,q\in Q_{T},E_{P,T}\rangle$. The models assign a probability to each triplet and select the highest probability question.
Haller RNN (Haller et al., 2022) An RNN-based model that orders word embeddings based on the fixation sequence and encodes each fixation using eye movement features. We adapt this model by appending question embeddings to the input and processing them through the RNN, followed by a classifier scoring each candidate question. Additional details on the model, including a diagram of the modified architecture and the used eye movement features, are presented in Section˜A.1.
RoBERTEye-Fixations (Shubi et al., 2024) A transformer-based model which integrates fixation-level eye movement features into RoBERTa (Liu et al., 2019). An eye movements feature vector for each fixation is first projected to the dimension of the word embeddings. These fixation vectors are then aligned with their corresponding words via the addition of position embeddings and finally concatenated with the word sequence. We adapt this model to incorporate candidate question embeddings and modify the classification layer to score questions. Section˜A.2 provides further details, including a model diagram and the eye movement features that the model uses.
@@ 4.2 Generative Models
We introduce two LLM-based approaches for generating the question based on the eye movements over the text. Both models are fine-tuned using an autoregressive next-token prediction objective for the correct question with a cross-entropy loss. During fine-tuning, the language models’ parameters remain frozen, and only an additional LoRA (Hu et al., 2022) adapter component is trained. As described in Section˜5, we evaluate these models both on the question selection and the question reconstruction tasks.
DalEye-LLaVA is based on LLaVA-OneVision (Li et al., 2024), a vision-language model that uses the Qwen 2-0.5B language model (Yang et al., 2024) as a backbone. The model input consists of a task prompt, the text, and eye movement feature embeddings for the text words. We replace the vision encoder of LLaVA-OneVision with a novel trainable eye movement encoder, which encodes eye movement features using fully connected and convolutional layers. The eye movements embeddings are positionally aligned with the text words. Additional details, including a model diagram and eye movement features, are provided in Section˜B.1.
DalEye-Llama In this approach, we provide the task and the eye movement trajectory of a participant over the text to the Llama 3.1 language model (Grattafiori et al., 2024) in textual form. The model is finetuned to reconstruct the true question given this input. We use a fixation-based eye movements representation, where each fixation is represented as a tuple containing the index of the fixated word, the word itself, the fixation duration in milliseconds, and the direction of the next saccade. This encoding preserves both the spatial locations and temporal order of the fixations on the text. It follows a prompt describing the task and instructing the model to generate the question that was presented to the reader. In Section˜B.2, we provide an example of the model input for a complete fixation sequence over a paragraph. We further report experimental results for a differently worded task prompt, as well as two alternative representations of the eye movement sequence, one based on word-level features and another combining word- and fixation-level features.
@@ 5.1 Data Splits
The OneStop eye tracking data is not i.i.d., each participant reads multiple paragraphs, and each paragraph is read by multiple participants. Following prior work on predictive modeling with OneStop (e.g. Shubi et al., 2024), we use a 10-fold cross-validation procedure where each of the 10 data splits has a training, validation, and a test set. To account for the structured nature of the data, the validation set and the test set are each partitioned into three disjoint parts that capture three different levels of model generalization, ordered by task difficulty:
New Participant: Training eye tracking data is available for the text but not for the participant.
@@
Each data split includes 64% of the trials in the training set, 17% in the validation set, and 19% in the test set – divided into 9% New Participant, 9% New Text, and 1% New Text & Participant. When aggregated across the 10 splits, 90% of the trials appear in both New Participant and New Text evaluation regimes, and 10% appear in the New Text & Participant regime. We ensure that in each split into training, validation, and test sets, there is a balanced distribution of question types, i.e. $q_{1,c_{1}},q_{2,c_{2}},q_{3,c_{2}}$ for each text in each part of the split. Hyperparameter optimization and model selection are conducted separately for each split. We assume that the evaluation regime with respect to novelty of items and participants at test time is unknown, and therefore model selection is performed using the entire validation set within each split. Further details on the hyperparameter search space, training, and evaluation procedures are provided in Appendix˜C.
5.2 Question Selection

### v3

# arXiv:2505.02872v3[cs.CL] 02 Mar 2026

@@ Abstract
When reading, we often have specific information that interests us in a text. For example, you might be reading this paper because you are curious about LLMs for eye movements in reading, the experimental design, or perhaps you wonder “This sounds like science fiction. Does it actually work?”. More broadly, in daily life, people approach texts with any number of text-specific goals that guide their reading behavior. In this work, we ask whether open-ended reading goals can be automatically decoded solely from eye movements in reading. To address this question, we introduce goal decoding tasks and evaluation frameworks using large-scale eye tracking for reading data in English with hundreds of text-specific information seeking tasks and auxiliary annotations of task-critical information. We develop and compare several discriminative and generative multimodal text and eye movements LLMs for these tasks. Our experiments show considerable success on selecting the correct goal among several options, and even progress towards free-form textual reconstruction of the precise goal formulation. We further tie model performance to cognitively interpretable aspects of human gaze behavior. These results open the door for further scientific investigation of goal driven reading, as well as the development of educational and assistive technologies that will rely on real-time decoding of reader goals from their eye movements. [^1] Code is available at https://github.com/lacclab/Open-Ended-Goal-Decoding.
@@ 1 Introduction
Eye movements in reading are a key methodology for studying the cognitive basis of reading and language processing. When we read, our eyes move over the text in a saccadic manner, with prolonged periods of time in which the gaze is stable, called fixations, and fast transitions between fixations called saccades (Schotter and Dillon, 2025). This trajectory contains rich behavioral traces of the ways in which readers interact with texts and process language in real time (Just and Carpenter, 1980; Rayner, 1998). Understanding the nature and strength of the relation between eye movements, the text, and online language processing has been a major research avenue in the psychology of reading in the past few decades, and in recent years has also been gaining increasing interest in NLP and machine learning (Barrett and Hollenstein, 2020; Reich et al., 2025).
In this work, we investigate a reading scenario that is prevalent in daily life but remains understudied – seeking specific information in a text. This scenario deviates from a common implicit assumption in psycholinguistics, according to which the comprehender’s goal is constant across communicative contexts: a general understanding of the linguistic input. In practice, readers approach texts with a variety of goals and information seeking needs. Each such goal can have a profound impact on the cognitive processes of language comprehension and the corresponding reading behavior (Radach and Kennedy, 2004; Kaakinen and Hyönä, 2010; Hahn and Keller, 2023; Shubi and Berzak, 2023).
@@
Tasks: We present a new cognitive state decoding challenge from eye movements in reading: decoding arbitrary information seeking goals over the content of specific texts. We instantiate this challenge as a goal selection task given several possible goals, and as an even more challenging goal reconstruction task where the information-seeking goal has to be generated.
Models: For the goal selection task, we introduce two strong baseline models and adjust two state-of-the-art discriminative models of text and eye movements (Haller RNN and RoBERTEye). For the goal reconstruction task, we develop and finetune two generative LLMs, with textual representations of eye movements (DalEye-Llama, DalEye-GPT). We further use textual eye movement representations to address this task using a frontier LLM (Gemini) in zero-shot and few-shot regimes.
Evaluations: We perform systematic evaluations of goal selection accuracy at two levels of task difficulty and utilize several generation quality measures for the goal reconstruction task. For both tasks, we examine model generalization across new texts and new readers. Finally, we tie model performance to cognitively interpretable properties of human gaze behavior during goal decoding.
@@ 2 Eye Tracking Data
We use OneStop (Berzak et al., 2025b), a public eye tracking dataset collected with an EyeLink 1000 Plus eye tracker. It comprises 360 adult native English speakers reading Guardian articles from the OneStopQA multiple choice reading comprehension dataset (Vajjala and Lučić, 2018; Berzak et al., 2020). In each experimental trial, a participant reads a paragraph on a screen and, on the following screen, answers a multiple-choice reading comprehension question, without the ability to return to the paragraph. Each paragraph can appear in its original or human-simplified form. 180 participants read in an information seeking condition, where they receive the question (but not the answers) before reading the paragraph, and therefore have a specific reading goal for the upcoming paragraph. The remaining 180 participants do not receive the question ahead of time, and thus have to be ready to answer any subsequent question. We use the information seeking part of OneStop (Berzak et al., 2025a), which is uniquely suited for our study.
For each paragraph, a participant receives one of three possible questions over the paragraph. Each such question has a manually annotated critical span in the paragraph, which contains the information that is essential for answering the question. The questions tend to be more inferential than extractive: typically, the critical span does not contain the answer in verbatim form, but rather the essential information from which it can be inferred (Berzak et al. (2020), and see Appendix˜A for an empirical analysis which supports this property). Two of the questions share the same critical span, while the third question has a distinct critical span. This design supports a question selection task with two tiers of difficulty, distinguishing between questions with non-overlapping critical spans and the more challenging variant of distinguishing between questions over the same critical span. Figure˜1 presents an example of a paragraph, its three questions, and their critical span annotations.
Figure 1: Example of a OneStop paragraph, its three questions, and each question’s critical span: the paragraph segment essential for answering the question. See details on notation in Section˜3.
OneStop Information Seeking has 486 questions over 162 paragraphs, with 20 participants answering each question. The mean question length is 10 words. The mean paragraph length is 109 words (5.8 sentences), of which the critical span is 32 words on average. Overall, there are 1,055,429 paragraph word tokens for which eye tracking data was collected.
@@ 3 Problem Formulation
The general problem we address is using eye movements over a text (in our setting, a text is always a paragraph) to decode which question a participant seeks to answer using that text. Each text $T$ has three text specific questions, $Q_{T}=\left\{q_{1,c_{1}},q_{2,c_{2}},q_{3,c_{2}}\right\}$. The critical information for answering question $q_{1,c_{1}}$ is located in the text in span $c_{1}$, while for both questions $q_{2,c_{2}}$ and $q_{3,c_{2}}$ it is in span $c_{2}$. In each experiment trial, a participant $P$ is presented with one question, $q_{P}\in Q_{T}$, before reading the text. The recording of the participant’s eye movements while reading the text is denoted by $E_{P,T}$. We propose the following two task variants.
Goal Selection In the goal selection task, a classifier $h$ has to select which question from $Q_{T}$ the participant received, from their eye movement recording over the text:
$h(T,E_{P,T},Q_{T})\longrightarrow\hat{q}_{P}\in Q_{T}$
Goal Reconstruction In the goal reconstruction task, a generative model $g$ has to generate the question that the participant received, from their eye movement recording over the text:
@@ 4 Models
Figure 2: Two model types for decoding the question presented to a participant before reading the text, from their eye movements over the text. (a) Discriminative models that score candidate questions and are evaluated on question selection accuracy. (b) Generative models that reconstruct the question presented to the reader, and are evaluated on question reconstruction quality.
@@ 4.1 Discriminative Models
We introduce two baseline models which build on the following observations from prior literature: (i) readers tend to spend more time on question-relevant than on question-irrelevant portions of the text (Kaakinen et al., 2002; Malmaud et al., 2020; Hahn and Keller, 2023; Shubi and Berzak, 2023), and (ii) question-relevant text segments tend to have higher semantic similarity to the question compared to question-irrelevant segments (Mitra et al., 2016).
@@
Question Similarity to RT-Weighted Passage (Figure˜3(a)) For a participant reading a text of $N$ words, we compute a text embedding, $\text{emb}(T)$, as a weighted average of its contextualized word embeddings, where the weight of each word is the fraction of time spent reading it. Formally: $\text{emb}(T)=\sum_{i=1}^{N}\text{RT}(w_{i})\,\text{emb}(w_{i})$. The embeddings for word $w_{i}$, $\text{emb}(w_{i})$, are extracted from RoBERTa (Liu et al., 2019), and $\text{RT}(w_{i})$ is the word’s Total Fixation Duration (i.e., sum of all the fixation durations) normalized by the text’s total reading time. For each candidate question $q_{j}$, we use its [CLS] token embedding as $\text{emb}(q_{j})$, and compute its cosine similarity with $\text{emb}(T)$. We then predict the question with the highest similarity: ${\arg\max}_{j\in\{1,_{\cdots},|Q_{T}|\}}\cos(\text{emb}(T),\text{emb}(q_{j}))$.
RT Similarity to Question-Word Similarities (Figure˜3(b)) An alternative to the model above, compares the vector of word reading times for the text to the vector of word–question similarities. Specifically, for a given candidate question $q_{j}\in Q_{T}$ and each text word $w_{i}$, we first compute $\text{sim}(q_{j},w_{i})=\cos\bigl(\text{emb}(q_{j}),\text{emb}(w_{i})\bigr)$, where as before, $\text{emb}(w_{i})$ is a contextual word embedding from RoBERTa, and $\text{emb}(q_{j})$ is the question [CLS] embedding, yielding a length-$N$ vector of similarities $\text{sim}(q_{j},T)=[\text{sim}(q_{j},w_{1}),\text{sim}(q_{j},w_{2}),...,\text{sim}(q_{j},w_{n})]$. We select the question maximizing the cosine similarity between this vector and $\text{RT}(T)$, the length-$N$ vector of speed-normalized Total Fixation Durations for the text words: ${\arg\max}_{j\in\{1,_{\cdots},|Q_{T}|\}}\cos\bigl(\text{sim}(q_{j},T),\,\text{RT}(T)\bigr)$.
@@
We adapt two state-of-the-art encoder models for eye movements and text. Building on the approach of Radford et al. (2019) for encoding multiple-choice QA items, we use several copies of each text, where instead of combining each copy with a candidate answer, we combine it with a candidate question, and further provide the participant’s eye movements over the text $\langle T,E_{P,T},q\in Q_{T}\rangle$. The models assign a probability to each triplet and select the highest probability question.
Haller RNN (Haller et al., 2022) An RNN model that represents each fixation on the text using eye movement features concatenated with the corresponding RoBERTa-based word embedding. We append an embedding of the question to the input of the model. The model scores each candidate question. Additional details on the model, including a diagram of the modified architecture and the used eye movement features, are presented in Section˜B.1.
RoBERTEye-Fixations (Shubi et al., 2024) A transformer-based model which integrates fixation-level eye movement features into RoBERTa. An eye movements feature vector for each fixation is first projected to the dimension of the word embeddings. These fixation vectors are then aligned with their corresponding words via the addition of position embeddings, and finally concatenated with the word sequence. We adapt this model to incorporate candidate question embeddings and modify the classification layer to score questions. Section˜B.2 provides further details, including a model diagram and the eye movement features that the model uses.
@@ 4.2 Generative Models
We experiment with two approaches for generating the question based on eye movements over the text. The first approach (DalEye-Llama, DalEye-GPT) is fine-tuning, where the task, the text, and the eye movements are provided to the model in the prompt in textual form. The second approach uses the same prompts for zero-shot and few-shot generation from a frontier LLM (Gemini).
DalEye-Llama We provide the task, the text, and the eye movement trajectory of a participant over the text to Llama 3.1 (Grattafiori et al., 2024) in textual form. The model is fine-tuned to reconstruct the true question given this input using an autoregressive next-token prediction objective. We use a fixation-based eye movements representation, where each fixation is represented as a tuple containing the index of the fixated word, the word itself, the fixation duration, and the direction of the next saccade. This encoding preserves both the spatial locations and the temporal order of the fixations. This input is preceded by the text and a task description instructing the model to generate the question that was presented to the reader. An example of a prompt is shown in Section˜C.1.
DalEye-GPT A similar approach to DalEye-Llama, with identical prompts, where instead we finetune a commercial LLM, GPT-4o-mini (Hurst et al., 2024).
Gemini We evaluate a frontier commercial LLM, Gemini-3-Pro-Preview (Google DeepMind, 2025) in zero-shot and few-shot settings. In the zero-shot setting, the model receives the same textual input as DalEye-Llama and DalEye-GPT. In the few-shot setting, the prompt further includes 10 randomly drawn trial examples with ground truth questions (see Section˜C.1.2 for details).
In Section˜C.2 we further provide details on DalEye-LLaVA, an alternative, embedding based approach to finetuning based on the LLaVA-OneVision language-vision model architecture (Li et al., 2025). Although this model mostly did not perform well on our task, we provide its description and evaluations in the Appendix as a reference point for future embedding based models.
@@ 5.1 Data Splits
Following prior work on predictive modeling with OneStop (e.g. Shubi et al., 2024), we use a 10-fold cross-validation procedure where each of the 10 data splits has a training, validation, and test set. The OneStop eye tracking data is not i.i.d.; each participant reads multiple paragraphs, and each paragraph is read by multiple participants. To account for this structured nature of the data, the validation set and the test set of each data split are each partitioned into three disjoint parts that capture three different levels of model generalization, ordered by task difficulty:
New Participant: Training eye tracking data is available for the text (with each of the three possible questions) but not for the participant.
@@
Each data split includes 64% of the trials in the training set, 17% in the validation set, and 19% in the test set – divided into 9% New Participant, 9% New Text, and 1% New Text & Participant. When aggregated across the 10 splits, 90% of the trials appear in both New Participant and New Text evaluation regimes, and 10% appear in the New Text & Participant regime. We ensure that in each split into training, validation, and test sets, there is a balanced distribution of question types, i.e. $q_{1,c_{1}},q_{2,c_{2}},q_{3,c_{2}}$ for each text in each part of the split. Hyperparameter optimization and model selection are performed separately for each split. We assume that the evaluation regime with respect to novelty of texts and participants at test time is unknown, and therefore model selection is performed using the entire validation set within each split. Further details on the hyperparameter search space, training, and evaluation procedures are provided in Appendix˜D.
5.2 Goal Selection